(*) If n is a positive integer such that sin180/2^n + cos 180/2^n = root of n / 2, then

(*)
If n is a positive integer such that sin180/2^n + cos 180/2^n = root of n / 2, then

Grade:11

1 Answers

Nishant Vora IIT Patna
9 years ago
Hello student, Please find the answer to your question below

sin180/2^n + cos 180/2^n = root of n / 2 …...(given in question )
squaring both sides
1 + 2 sin180/2^n * cos 180/2^n = n/4

1 + Sin (180/ 2 n-1) = n/4
Now you can use hit and trial to get the ans